Farm Fresh Sdn Bhd supplies fresh chicken meat to retailers located throughout Petaling Jaya. The market or fresh chicken meat is perfectly competitive and the company estimated its total cost to be as follow:
〖TC=125+2Q+0.002Q〗^2
Calculate Farm Fresh Sdn Bhd’s optimal output and profit if fresh chicken meat prices are stable at RM4 a kilogram.
Calculate the industry price necessary for the firm to supply 1000 kilograms of fresh chicken meat.
1
Expert's answer
2020-07-07T10:18:46-0400

In equilibrium P = MC = MR.

"MC = TC'(Q) = 2 + 0.004Q."

If Q = 1,000, then P = MC = 2 + 0.004×1,000 = 6.